OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (2, 2) (3, 3) Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (2, 2) (3, 3) @Mathematics

OpenStudy (anonymous):

y=x

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (1, -5) (5, -1)? do you knbow this one

OpenStudy (anonymous):

y=x-6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)